OpenStudy (anonymous):

12=(m^2-5m)^2+(m^2-5m) find m

OpenStudy (anonymous):

\[12=(m ^{2}-5m)^2+(m ^{2}-5m)\]

OpenStudy (anonymous):

12=(m^2-5m)^2+(m^2-5m) 12 = m(m^3 - 10m^2 + 26x - 5) x = -.5413, 1, 4, 5.5 -- used graphing calcu

OpenStudy (anonymous):

I need to know how to do it algebraically

OpenStudy (anonymous):

nevermind. I solved it. thanks anyways